OpenStudy (anonymous):

For the parent function y√x what effect does a value of k = -3 have on the graph? Thanks!! :)

OpenStudy (jdoe0001):

\(y = \sqrt{3}\) ?

OpenStudy (anonymous):

Whoops forgot to put the multiple choice answers: -Horizontal shift of 3 units to the right -Horizontal shift of 3 units to the left -Vertical shift of 3 units up -Vertical shift of 3 units down

OpenStudy (jdoe0001):

woops, darnh I meant => \(y = \sqrt{x}\) ?

OpenStudy (anonymous):

Uhhh I'm obviously a retard in math so could you use actual words please lmfao sorry xD

OpenStudy (jdoe0001):

hmmm, well you see I dunno what meant , I mean... is -3 "ADDED" to it or part of the root or a multiplier of it?

OpenStudy (jdoe0001):

\(\bf k\sqrt{x}, \sqrt{x}-k, \sqrt{x-k} \ \ ?\)